A certain weak acid, HA, has a Ka value of 5.3×10−7

Calculate the percent ionization of HA in a 0.10 M solution.
Answer #1

HA dissociates as:

HA -----> H+ + A-

0.1 0 0

0.1-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((5.3*10^-7)*0.1) = 2.302*10^-4

since c is much greater than x, our assumption is correct

so, x = 2.302*10^-4 M

% dissociation = (x*100)/c

= 2.302*10^-4*100/0.1

= 0.2302 %

Answer: 0.23 %